# A comparison between humans and AI at recognizing objects in unusual poses

### v2

@@ Abstract
Deep learning is closing the gap with human vision on several object recognition benchmarks. Here we investigate this gap for challenging images where objects are seen in unusual poses. We find that humans excel at recognizing objects in such poses. In contrast, state-of-the-art deep networks for vision (EfficientNet, SWAG, ViT, SWIN, BEiT, ConvNext) and state-of-the-art large vision-language models (Claude 3.5, Gemini 1.5, GPT-4) are systematically brittle on unusual poses, with the exception of Gemini showing excellent robustness in that condition. As we limit image exposure time, human performance degrades to the level of deep networks, suggesting that additional mental processes (requiring additional time) are necessary to identify objects in unusual poses. An analysis of error patterns of humans vs. networks reveals that even time-limited humans are dissimilar to feed-forward deep networks. In conclusion, our comparison reveals that humans and deep networks rely on different mechanisms for recognizing objects in unusual poses. Understanding the nature of the mental processes taking place during extra viewing time may be key to reproduce the robustness of human vision in silico.
Machine Learning, ICML
@@ 1 Introduction
However, studies that investigate out-of-distribution generalization mostly focus on local distortions (e.g., texture modifications, blur, color modifications) as opposed to transformations that affect the global structure of the image, such as a change in viewpoint. A few studies have investigated the generalization capability of deep networks to recognize objects in various poses, both in non-adversarial (Alcorn et al., 2019; Ibrahim et al., 2022; Abbas & Deny, 2023; Madan et al., 2022) and adversarial settings (Zeng et al., 2019; Madan et al., 2021), and they show a substantial degradation of network performance for unusual poses. However, a direct comparison with human vision has been missing from these studies, leaving open the question of whether humans outperform networks on this task.
@@ 2.1 Dataset collection
We chose 51 different object categories from the ImageNet classes (see Appendix C for the list of objects), and obtained a corresponding 3D model for each category from Sketchfab (https://sketchfab.com/). The objects were chosen based on a few criteria:
@@
The stimulus set applied here is not entirely ecologically realistic, as all objects are displayed without a background. However, our aim was to create a highly controlled stimulus set and experimental setup. Performance in canonical view is compared with performance with exactly the same object in a rotated view, both shown without background. Thus, any performance differences, be it human or network, can only be caused by rotation. Conversely, any robustness can only be caused by rotation-invariance, not context-based reasoning.
@@ 2.2.1 Observers
Altogether 24 observers (12 women, 12 men, aged 19-54) participated in the experiments. Our observers comprised Finnish students from diverse academic backgrounds, including disciplines such as engineering, psychology, business, and medicine. Participants were required to have normal vision or corrected-to-normal vision. Additionally, the observers did not report any medical condition (epilepsy, migraine, etc.) that could affect the results.
@@ 2.3 Machine tests
We implemented the same tests on a collection of 5 state-of-the-art deep networks for vision (referred below as ’pure vision networks’) and 6 large vision-language models (VLM). Since EfficientNet was used to select the problematic images and labels, we did not include EfficientNet in this collection nor in our study of network robustness. We, in fact, investigated the transfer of errors between EfficientNet and these other networks.
We tested the 5 following pure vision networks: SWAG-RegNetY (Singh et al., 2022), BEiT-L/16 (Bao et al., 2021), ConvNext-XL (Liu et al., 2022), SWIN-L (Liu et al., 2021), and ViT-L/16 (Dosovitskiy et al., 2020) (see Appendix B for model descriptions). The network selection was based on previous work from Abbas & Deny (2023), which thoroughly evaluated the robustness of 37 state-of-the-art networks to objects presented in unusual poses. This collection provides a diverse set of state-of-the-art and very large models trained on large datasets (e.g., Imagenet-21K, Instagram 3.6B) with various architectures (ConvNet and Transformers) and training objectives (supervised and self-supervised). These networks obtained some of the most robust results to unusual poses according to Abbas & Deny (2023), and all of them were fine-tuned to the categories of ImageNet.
@@
Our selection of VLMs included Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o and GPT-4-vision-preview. The experiment was conducted via the API. Each model was shown the 147 images and provided with the following prompt (see examples in Appendix A):
@@ 2.4 Statistical tests
Figure 3: Comparing neural networks and humans at recognizing objects in various poses. Dark grey bars: Average performance of pure vision deep networks (left: upright vs. right: rotated) with 95% confidence intervals (n=5, grey points represent individual network performances). Light grey bars: Average performance of six large vision-language models (VLMs). Diamonds indicate individual model performances (full diamonds show the best-performing version of the different model classes). Blue bars: Average human performance with unlimited viewing time (n=12, grey points represent individual performances). Orange & Red bars: Average human performance with limited viewing time (150 ms and 40 ms, respectively, n=12, grey points represent individual performances). Chance performance is 50%. Three stars (***) indicate highly significant differences (p$<$0.001), ”n.s.” for not significant. With unlimited time, humans excel at recognizing rotated objects, while pure vision networks struggle (best: SWAG at 70.1%). Both GPT-4 and Claude models follow the same pattern as the pure vision networks, with a significant drop in accuracy for rotated images compared to upright ones. However, Gemini 1.5 Flash mirrors human performance with unlimited viewing time, achieving 97.7% accuracy on rotated images compared to human accuracy of 98.9%. Limiting human viewing time (40 ms or 150 ms) impairs their ability to recognize rotated objects, substantially more than upright objects, bringing their performance closer to the network level.
@@
where $p_{i}p_{j}$ is the probability that observes give the same correct response by chance and $(1-p_{i})(1-p_{j})$ the probability that they give the same incorrect response by chance. To evaluate the consistency between the two observers i and j, and determine whether it goes beyond what could be expected by chance, we use a metric called error consistency (Geirhos et al., 2020, 2021) (i.e., Cohen’s $\kappa$) given by:
@@ 3 Results
We compared human and machine vision on a task consisting in recognizing objects in unusual poses. For this, we selected 5 different pure vision deep neural networks (SWAG, ViT, SWIN, BEiT, ConvNext), chosen for their state-of-the-art performance in recognizing objects in unusual poses (Abbas & Deny, 2023), and 6 state-of-the-art large vision-language models (Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o, GPT-4-vision-preview). We performed a two-alternative-forced-choice object categorization task: an image of an object was presented to the viewer (either deep network or human participant), and then the viewer had to choose between two different names (i.e., labels) for the object. The two labels were selected based on the highest activations of Noisy Student EfficientNet’s output layer (Xie et al., 2020), the most robust network to unusual poses according to Abbas & Deny (2023). For pure vision networks, the choice was made by looking at the highest activation of the softmax output layer for these two labels. EfficientNet was excluded from our analyses comparing humans and networks because images and labels were chosen based on its mistakes (see Methods). For the vision-language models, each VLM was given an image and prompted with the 2 label choices, and the label chosen by the model was selected to be its answer. The performance of VLMs is discussed in the last paragraph of results, while previous paragraphs concentrate on the comparison of pure vision networks and humans.
@@
Figure 4: Error patterns are different for neural networks and time-limited humans. a) Comparison of human and network accuracy for the rotated-correct condition (EfficientNet was correct on these rotations) vs. rotated-incorrect condition (rotations that have failed EfficientNet). Humans show consistent accuracy between rotated-correct and rotated-incorrect conditions. In contrast, networks, including the six VLMs (represented by diamonds), exhibit a performance drop. However, for Gemini 1.5 Flash, Gemini 1.5 Pro and Claude Opus, the decrease in performance between the two conditions is not as notable. b) Error consistency analysis performed on the 5 neural networks (not including VLMs) and 40 ms time-limited human subjects. 12 human subjects were partitioned into 4 groups of 3, so that every group saw each rotated image exactly once. The darker red cluster for networks (dark red = highly consistent errors) indicates that they have similar patterns of error, which are not shared by human subjects, highlighting that EfficientNet errors transfer better to other networks than to humans. c) Mean error consistencies were calculated by comparing networks with each other and with human subjects (i.e., average computed over the different matrix clusters). A two-tailed unpaired t-test confirms that networks indeed make more consistent errors with each other than with humans (t(28) = 4.0, p = 3.8e-04).
@@
We found that—even under acute time limitation—humans rely on different strategies than networks for recognition, relying more on object structure than their network counterparts (which rely on object details). In Figure 5a, we show examples where both networks and humans performed well. In these examples, both the structure and details of the objects are well informative about the identity of the object. In Figure 5b, we show examples of errors that networks made but humans typically didn’t. A striking example is the image of a stretcher, which networks mislabeled as a matchstick. We speculate that the similarity between the stretcher’s legs and matchsticks is the cause for this misclassification. Despite the resemblance to matchsticks, humans however correctly identified the object because the overall structure of the object rules out the possibility of matchsticks. Thus, we can reason that in that case, humans tended to perceive the overall structure rather than fixating on minute details, unlike networks which based their prediction on such details. In Figure 5c, we show examples where identification of an object is challenging for humans but not for networks. For instance, images of jellyfish posed difficulties for human observers. The shortness of the exposure time, 40 ms, could potentially limit humans to recognising only the overall shape and color of the object, compatible with both a lampshade and a jellyfish. In contrast, networks are able to process fine details, such as the patterns on the jellyfish’s umbrella, allowing for accurate identification regardless of pose. Figure 5d presents examples that prove to be challenging for both humans and networks. It can be noted that in all these examples, both the structure and details contain only limited information about the object identity. The impairment of both networks and humans on these images is thus consistent with our hypothesis that humans rely on structure and networks on details for recognition. These qualitative observations reinforce our conclusion that time-limited humans are not well-modelled by feed-forward networks.
@@
Next, we tested six large vision-language models (VLMs): GPT-4o, GPT-4-vision-preview, Claude 3.5 Sonnet, Claude 3 Opus, Gemini 1.5 Flash and Gemini 1.5 Pro.
@@
With upright poses, all 6 models performed well (Gemini 1.5 Flash, GPT-4o and GPT-4-vision-preview achieving 100% accuracy, Gemini 1.5 Pro 98.0%, Claude 3.5 Sonnet 94.1% and Claude 3 Opus 92.2.%). However, there were significant differences in performance when recognizing rotated poses (Figure 3). Claude 3 Opus performed below the 50% chance level, achieving only 45.8% accuracy. Claude 3.5 Sonnet’s accuracy was on par with pure vision networks with 62.5% accuracy. GPT-4o had an accuracy of 84.8% and GPT-4-vision-preview 79.7%, performing substantially better than our best performing pure vision model, SWAG (70.1%). Remarkably, Gemini 1.5 Pro achieved an accuracy of 96.6% and Gemini 1.5 Flash an accuracy of 97.7.%, performances on par with unlimited humans (98.9%). Separating the rotated condition into two categories—rotated-correct (images correctly labelled by EfficientNet) and rotated-incorrect (images incorrectly labelled by EfficientNet)—revealed that the 6 VLMs exhibited a performance drop between these conditions, similarly to the five pure vision networks (Figure 4a). Performance differences of GPT-4o (18%), GPT-4-vision-preview (17%) and Claude 3.5 Sonnet (32%) were the most similar to the 5 pure vision networks, whose average performance gap was 26%. For Gemini models and Claude 3 Opus, the performance differences were much smaller: 4.7% for Gemini 1.5 Flash, 2.7% for Gemini 1.5 Pro and 2.4% for Claude. For both unlimited and time-limited humans, the performance difference was not significant between the two conditions. This suggests that errors from EfficentNet transfer better to most VLMs than to humans.
@@ 4 Discussion
Our results strikingly demonstrate that the human visual system is more robust than deep networks for vision and than most vision-language models at recognizing objects in unusual poses. Indeed, with unlimited viewing time, human performance is not affected by unusual poses, whereas deep networks are fooled by them. Only in the very limited viewing time condition (40 ms) does the performance of humans deteriorate to the level of deep networks.

### v3

@@ Abstract
Deep learning is closing the gap with human vision on several object recognition benchmarks. Here we investigate this gap in the context of challenging images where objects are seen in unusual poses. We find that humans excel at recognizing objects in such poses. In contrast, state-of-the-art deep networks for vision (EfficientNet, SWAG, ViT, SWIN, BEiT, ConvNext) and state-of-the-art large vision-language models (Claude 3.5, Gemini 1.5, GPT-4, SigLIP) are systematically brittle on unusual poses, with the exception of Gemini showing excellent robustness to that condition. As we limit image exposure time, human performance degrades to the level of deep networks, suggesting that additional mental processes (requiring additional time) are necessary to identify objects in unusual poses. An analysis of error patterns of humans vs. networks reveals that even time-limited humans are dissimilar to feed-forward deep networks. In conclusion, our comparison reveals that humans are overall more robust than deep networks and that they rely on different mechanisms for recognizing objects in unusual poses. Understanding the nature of the mental processes taking place during extra viewing time may be key to reproduce the robustness of human vision in silico. All code and data is available at https://github.com/BRAIN-Aalto/unusual_poses.equal contribution as last authors.
@@ 1 Introduction
However, studies that investigate out-of-distribution generalization mostly focus on local distortions (e.g., texture modifications, blur, color modifications) as opposed to transformations that affect the global structure of the image, such as a change in viewpoint. A few studies have investigated the generalization capability of deep networks to recognize objects in various poses, both in non-adversarial (Alcorn et al., 2019; Ibrahim et al., 2022; Madan et al., 2022; Abbas & Deny, 2023) and adversarial settings (Zeng et al., 2019; Madan et al., 2021b; a), and they show a substantial degradation of network performance for unusual poses. However, a direct comparison with human vision has been missing from these studies, leaving open the question of whether humans outperform networks on this task.
@@ 2.1 Dataset collection
We chose 51 different object categories from the ImageNet classes (see Appendix D for the list of objects), and obtained a corresponding 3D model for each category from Sketchfab (https://sketchfab.com/). The objects were chosen based on a few criteria:
@@
Caveat 1:
We used EfficientNet to identify challenging object poses and label alternatives. A caveat of this selection method is that it may introduce a bias against networks, since half of the images were selected to fool EfficientNet (which is a neural network). To mitigate this bias, we separately compare networks and humans on the half of images that have not failed EfficientNet (rotated-correct condition in Fig. 4). Essentially, our conclusions remain identical on that subset: humans (nearly 100% accurate) are much more robust than pure-vision deep networks (overall best network of our collection: SWAG: 82% accuracy) and of most vision-language models (except Gemini). Additionally, we note that our selection is hardly adversarial, since EfficientNet failed on a large fraction of random poses (14.5% of poses, cf Abbas & Deny (2023)). Our study is thus not akin to studies performing fine pose optimization to find adversarial examples (e.g., Madan et al. (2021a)).
Caveat 2:
The stimulus set applied here is not entirely ecologically realistic, as all objects are synthetic 3D models and are displayed without a background. However, these choices are justified by our objective to create a highly controlled stimulus set and experimental setup: performance in canonical view is compared with performance in a rotated view of the very same object, both shown without background. Thus, any performance differences between these conditions, be it human or network, can only be caused by rotation. Conversely, any robustness found across conditions can only be due to rotation-invariant recognition, not context-based reasoning.
@@ 2.2.1 Observers
Altogether 24 observers (12 women, 12 men, aged 19-54) participated in the experiments. Our observers comprised Finnish students from diverse academic backgrounds, including disciplines such as engineering, psychology, business, and medicine. Participants were required to have normal vision (i.e., individuals using vision aids such as glasses to achieve normal vision) or corrected-to-normal vision. Additionally, the observers did not report any medical condition (epilepsy, migraine, etc.) that could affect the results.
@@ 2.3 Machine tests
We then implemented the same tests on a collection of 5 state-of-the-art deep networks for vision (referred below as ’pure vision networks’) and 6 large vision-language models (VLM). Since EfficientNet was used to select the problematic images and labels, we did not include EfficientNet in this collection nor in our study of network robustness.
We tested the 5 following pure vision networks: SWAG-RegNetY (Singh et al., 2022), BEiT-L/16 (Bao et al., 2021), ConvNext-XL (Liu et al., 2022), SWIN-L (Liu et al., 2021), and ViT-L/16 (Dosovitskiy et al., 2020) (see Appendix C for model descriptions). The network selection was based on previous work from Abbas & Deny (2023), which thoroughly evaluated the robustness of 37 state-of-the-art networks to objects presented in unusual poses. This collection provides a diverse set of state-of-the-art and very large models trained on large datasets (e.g., Imagenet-21K, Instagram 3.6B) with various architectures (ConvNet and Transformers) and training objectives (supervised and self-supervised). These networks obtained some of the most robust results to unusual poses according to Abbas & Deny (2023), and all of them were fine-tuned to the categories of ImageNet.
@@
Our selection of VLMs included Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o, GPT-4-vision-preview and SigLIP (Zhai et al., 2023). For the large-language models (all models excluding SigLIP), the experiment was conducted via the API. Each model was shown the 147 images and provided with the following prompt (see examples in Appendix A):
@@
For SigLIP, a state-of-the-art contrastive text-image model (successor of CLIP (Radford et al., 2021)), we fed it the image as well as the two label options. The label with the highest score of similarity was accepted as the answer of the model.
@@ 2.4 Statistical tests
Figure 3: Comparing neural networks and humans at recognizing objects in various poses. Dark grey bars: Average performance of pure vision deep networks (left: upright vs. right: rotated) with 95% confidence intervals (n=5, grey points represent individual network performances). Light grey bars: Average performance of seven large vision-language models (VLMs). Diamonds indicate individual model performances (full diamonds show the best-performing version of the different model classes). Blue bars: Average human performance with unlimited viewing time (n=12, grey points represent individual performances). Orange & Red bars: Average human performance with limited viewing time (150 ms and 40 ms, respectively, n=12, grey points represent individual performances). Chance performance is 50%. Three stars (***) indicate highly significant differences (p$<$0.001), "n.s." for not significant. With unlimited time, humans excel at recognizing rotated objects, while pure vision networks struggle (best: SWAG at 70.1%). GPT-4, Claude and SigLIP models follow the same pattern as the pure vision networks, with a significant drop in accuracy for rotated images compared to upright ones. However, Gemini 1.5 Flash mirrors human performance with unlimited viewing time, achieving 97.7% accuracy on rotated images compared to human accuracy of 98.9%. Limiting human viewing time (40 ms or 150 ms) impairs their ability to recognize rotated objects, substantially more than upright objects, bringing their performance closer to network levels.
@@
where $p_{i}p_{j}$ is the probability that observes give the same correct response by chance and $(1-p_{i})(1-p_{j})$ the probability that they give the same incorrect response by chance. To evaluate the consistency between the two observers i and j, and determine whether it goes beyond what could be expected by chance, we use a metric called error consistency (Geirhos et al., 2020; 2021) (i.e., Cohen’s $\kappa$) given by:
@@ 3 Results
We compared human and machine vision on a task consisting in recognizing objects in unusual poses. For this, we selected 5 different pure vision deep neural networks (SWAG, ViT, SWIN, BEiT, ConvNext), chosen for their state-of-the-art performance in recognizing objects in unusual poses (Abbas & Deny, 2023), and 7 state-of-the-art large vision-language models (Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o, GPT-4-vision-preview, SigLIP). We performed a two-alternative-forced-choice object categorization task: an image of an object was presented to the viewer (either deep network or human participant), and then the viewer had to choose between two different names (i.e., labels) for the object. The two labels were selected based on the highest activations of Noisy Student EfficientNet’s output layer (Xie et al., 2020), the most robust network to unusual poses according to Abbas & Deny (2023). For pure vision networks, the choice was made by looking at the highest activation of the softmax output layer for these two labels. EfficientNet was excluded from our analyses comparing humans and networks because images and labels were chosen based on its mistakes (see Methods). For the vision-language models, each VLM was given an image and prompted with the 2 label choices, and the label chosen by the model was selected to be its answer. The performance of VLMs is discussed in the last paragraph of results, while previous paragraphs concentrate on the comparison of pure vision networks and humans.
@@
Figure 4: Error patterns are different for neural networks and time-limited humans. a) Comparison of human and network accuracy for the rotated-correct condition (EfficientNet was correct on these rotations) vs. rotated-incorrect condition (rotations that have failed EfficientNet). Humans show consistent accuracy between rotated-correct and rotated-incorrect conditions. In contrast, networks, including the seven VLMs (represented by diamonds), exhibit a performance drop. However, for Gemini 1.5 Flash, Gemini 1.5 Pro and Claude Opus, the decrease in performance between the two conditions is not as notable. b) Error consistency analysis performed on the 5 neural networks (not including VLMs) and 40 ms time-limited human subjects. 12 human subjects were partitioned into 4 groups of 3, so that every group saw each rotated image exactly once. The darker red cluster for networks (dark red = highly consistent errors) indicates that they have similar patterns of error, which are not shared by human subjects, highlighting that EfficientNet errors transfer better to other networks than to humans. c) Mean error consistencies were calculated by comparing networks with each other and with human subjects (i.e., average computed over the different matrix clusters). A two-tailed unpaired t-test confirms that networks indeed make more consistent errors with each other than with humans (t(28) = 4.0, p = 3.8e-04).
@@
We found that—even under acute time limitation—humans rely on different strategies than networks for recognition, relying more on object structure than their network counterparts (which rely on object details). In Figure 5a, we show examples where both networks and humans performed well. In these examples, both the structure and details of the objects are well informative about the identity of the object. In Figure 5b, we show examples of errors that networks made but humans typically didn’t. A striking example is the image of a stretcher, which networks mislabeled as a matchstick. We speculate that the similarity between the stretcher’s legs and matchsticks is the cause for this misclassification. Despite the resemblance to matchsticks, humans however correctly identified the object because the overall structure of the object rules out the possibility of matchsticks. Thus, we can reason that in that case, humans tended to perceive the overall structure rather than fixating on minute details, unlike networks which based their prediction on such details. In Figure 5c, we show examples where identification of an object is challenging for humans but not for networks. For instance, images of jellyfish posed difficulties for human observers. The shortness of the exposure time, 40 ms, could potentially limit humans to recognising only the overall shape and color of the object, compatible with both a lampshade and a jellyfish. In contrast, networks are able to process fine details, such as the patterns on the jellyfish’s umbrella, allowing for accurate identification regardless of pose. Figure 5d presents examples that prove to be challenging for both humans and networks. It can be noted that in all these examples, both the structure and details contain only limited information about the object identity. The impairment of both networks and humans on these images is thus consistent with our hypothesis that humans rely on structure and networks on details for recognition. These qualitative observations reinforce our conclusion that time-limited humans are not well-modelled by feed-forward networks. More examples of time-limited human failures are shown in App. B.
@@
Next, we tested seven large vision-language models (VLMs): GPT-4o, GPT-4-vision-preview, Claude 3.5 Sonnet, Claude 3 Opus, Gemini 1.5 Flash, Gemini 1.5 Pro and SigLIP.
@@
With upright poses, all 6 models performed well (Gemini 1.5 Flash, GPT-4o and GPT-4-vision-preview achieving 100% accuracy, Gemini 1.5 Pro and SigLIP 98.0%, Claude 3.5 Sonnet 94.1% and Claude 3 Opus 92.2.%). However, there were significant differences in performance when recognizing rotated poses (Figure 3). Claude 3 Opus performed below the 50% chance level, achieving only 45.8% accuracy. Claude 3.5 Sonnet’s accuracy was on par with pure vision networks with 62.5% accuracy. GPT-4o had an accuracy of 84.8%, SigLIP 84.0% and GPT-4-vision-preview 79.7%, performing substantially better than our best performing pure vision model, SWAG (70.1%). Remarkably, Gemini 1.5 Pro achieved an accuracy of 96.6% and Gemini 1.5 Flash an accuracy of 97.7.%, performances on par with unlimited humans (98.9%). Separating the rotated condition into two categories—rotated-correct (images correctly labelled by EfficientNet) and rotated-incorrect (images incorrectly labelled by EfficientNet)—revealed that the 7 VLMs exhibited a performance drop between these conditions, similarly to the five pure vision networks (Figure 4a). Performance differences of GPT-4o (18%), GPT-4-vision-preview (17%) and Claude 3.5 Sonnet (32%) were the most similar to the 5 pure vision networks, whose average performance gap was 26%. For Gemini models and Claude 3 Opus, the performance differences were much smaller: 4.7% for Gemini 1.5 Flash, 2.7% for Gemini 1.5 Pro and 2.4% for Claude 3 Opus. For both unlimited and time-limited humans, the performance difference was not significant between the two conditions. This suggests that errors from EfficentNet transfer better to most VLMs than to humans.
@@ 4 Discussion
Our results demonstrate that the human visual system is more robust than deep networks for vision and than most vision-language models at recognizing objects in unusual poses. Indeed, with unlimited viewing time, human performance is not affected by unusual poses, whereas deep networks are fooled by them. Only in the very limited viewing time condition (40 ms) does the performance of humans deteriorate to the level of deep networks.